### v2

# The DBL Survey I: discovery of 34 double-lined double white dwarf binaries

Abstract
[…]
keywords:
binaries: spectroscopic – stars: white dwarfs – supernovae: general1 Introduction
[…]
The first definitive discovery of a compact DWD binary was L8702, identified as a DA+DA and double-lined system with an orbital period of 1.6 d (Saffer et al., 1988), sparking the inspiration for dedicated searches in the decade that followed (Robinson & Shafter, 1987; Bragaglia et al., 1990; Foss et al., 1991; Marsh et al., 1995; Marsh, 1995; Moran et al., 1997). Since then and predominantly in the last two decades, there has been a boom in the number of compact DWD binaries discovered. Many stem from the pioneering work of two surveys: the ESO supernovae type Ia progenitor survey (SPY) (Napiwotzki et al., 2020) and the Extremely Low Mass (ELM) survey (Brown et al., 2020; Kosakowski et al., 2023). Of the 643 DA WDs with measured RVs observed within the SPY survey, 39 double degenerate binaries were identified; 20 being double-lined and 19 showing single-lined RV variability (Napiwotzki et al., 2001, 2002; Karl et al., 2003; Nelemans et al., 2005; Geier et al., 2010, 2011; Napiwotzki et al., 2020). From the ELM survey, Brown et al. (2020) report a final northern sample of 98 DWD binaries and thusfar a further 34 ELM WD binaries have been found in the southern sky (Kosakowski et al., 2020, 2023). Considering that the current sample size of compact DWD binaries comprises approximately 270 systems, together these surveys constitute the discovery of half of the sample, highlighting the need to be careful when arriving at conclusions about the population. Both surveys have clear observational biases that, while possible to include, can be difficult to incorporate into model comparisons (Kilic et al., 2011; Li et al., 2019). For instance, the SPY sample relied on white dwarf catalogues built from surveys targeting blue objects; as a result, it is biased against reddened regions and against cooler white dwarfs. The ELM survey has well defined colour (and hence temperature) selection criteria, however there is a priority in follow-up campaigns for the shorter orbital periods and their selection on orbital periods is less documented.
Observational biases also affect photometric surveys, which have been the leading means of identifying short-period double-degenerate binaries through their photometric variability (e.g. Burdge et al., 2020; Keller et al., 2022; van Roestel et al., 2022; Ren et al., 2023). Such systems probe the environment of DWDs on the cusp of merger, detonation or the survival of the binary. Yet naturally, photometric searches are more sensitive to stars that have larger radii (therefore lower WD masses) and are hotter due to a combination of reflection effects, eclipse likelihood and the improved photometric precision of bright systems. In the context of DWDs, this leads to an unavoidable deficit of high-mass binaries in photometric searches and the only means to overcome large observational biases is through complete, volume-limited samples (Toonen et al., 2017).
The drawback is that volume-limited samples depend on a smaller subset of systems since complete samples out to large distances are observationally expensive. Recently, all-sky, multi-object spectrograph surveys have provided a means to improve the statistics for synthetic population comparison due to the sheer number of WDs of all types that can be observed (Badenes & Maoz, 2012). Through a combined comparison with the Sloan Digital Sky Survey (SDSS) and the SPY survey, Maoz et al. (2018) speculate that 10% of the single WD population are a result of DWDs merger events, serving as a means to explain a ‘broad shoulder’ of WDs between 0.7–0.9 (see also Tremblay et al., 2016; Cunningham et al., 2024, for other interpretations on the shoulder in the DA mass distribution). Meanwhile, Kilic et al. (2020b) find that mergers of all kinds (Temmink et al., 2020) are a contributing factor but are insufficient to reproduce the broad shoulder in volume-limited populations.
A new means of detecting thousands of wide binary companions to a WD has lately become possible through astrometric solutions using Gaia (Gaia Collaboration et al., 2016, 2023). El-Badry & Rix (2018); El-Badry et al. (2021) revealed hundreds of wide DWD binaries through the means of a common proper motion and parallax between the two stars (100 au). As for the more compact systems, signatures of binarity in the local population can be predicted through a low-amplitude astrometric wobble of the centre of light, which has often been used to select strong candidates through the so called Renormalised Unit Weight Error (RUWE, Belokurov et al., 2020). Of the likes, Korol et al. (2022b) analyse RUWE-excessive candidates to identify a gap in the separation distribution of DWDs at approximately 1 au while finding a consistent DWD fraction with the spectroscopic sample analyses and indicating a sensitivity in the detection of DWDs with separations of 0.01 au < a < 2 au.
Looking towards the future, upcoming Gaia data releases should begin to expand the time baseline of astrometric measurements to probe DWDs separations <0.01 au and reveal many more DWD candidates at further distances through time-series photometry (Steen et al., 2024). An ever growing completeness in the spectroscopically observed sample will build better sensitivity for radial velocity (RV) variations over all orbital periods, and the gravitational-wave signature of ultra-compact binaries across the entire Milky Way will be unlocked through the launch of the Laser Interferometer Space Antenna (LISA, Amaro-Seoane et al., 2023b), providing a few thousands of systems (Korol et al., 2017, 2018, 2022a; Amaro-Seoane et al., 2023a; Kupfer et al., 2024). However, while each of these methods have great potential to interpret the orbital parameters of the binary alone, separating one star from the other is far more challenging and relies on a unique signature from both of the stars in the data, best identifiable through eclipsing systems and/or spectroscopically double-lined DWDs.
Of the full confirmed and compact DWD sample, where we have created and made accessible a database that will be regularly maintained at github.com/JamesMunday98/CloseDWDbinaries, just 46 are known to be double-lined and 20 have a mass precision for both stars of . Only with a large sample of precise constraints for both stars can we unravel the nuances of binary star evolution and effectively construct evolutionary models suitable for all initial conditions of two stars. In this pilot study, we introduce the double-lined double white dwarf (DBL) survey which aims to expand the double-lined and well-constrained sample of compact DWDs significantly, constructed with clearly defined selection cuts based on the position of candidate systems in the Hertzsprung-Russell (HR) diagram. As well, we strive to provide a well-constrained sample of compact DWDs surveying a range of orbital separations/periods to facilitate better mapping of post-common envelope binary stars. The undertaken search indicates an approximately one in three success rate of a candidate being double-lined, outlining a clear means to expand to an observable population of precisely constrained DWD binaries and especially so with the addition of phase-resolved spectroscopy (Munday et al, in prep).
2 The double-lined double white dwarf survey sample
2.1 Sample selection
Beginning the survey in 2018, we initiated the target selection of the DBL survey using the latest Gaia data release available at the time: Gaia DR2 (Gaia Collaboration et al., 2018). We started with the full catalogue of 486 641 candidate WD sources presented in Gentile Fusillo et al. (2019), all of which assigned a probability of being a WD (Pwd).
In search of double-lined DWD binaries, we considered a few steps. Firstly, we consider that the majority of the population of WDs fall along a typical cooling track of a 0.6 DA WD and use this as a baseline for the expected luminosity of a single star. As shown in Fig. 1, this track traverses the A-branch of WDs that is dominated by a population of hydrogen-rich atmosphere DA WDs (Jiménez-Esteban et al., 2023), implying that the majority of systems were expected to have Balmer absorption lines usable for RV extraction. Considering that we sought to find double-lined DWDs where the dimmer companion is not washed out of the spectrum by the brighter star, we aimed for the temperature of the brighter star to be less than 20 000 K. Such a 0.6 WD has mag, therefore, all targets in the selection required mag.
Next, we required that the flux ratio between the companion and the brighter star is greater than a third to show a clear spectroscopic signature, meaning that a second star would cause the binary to be 0.31 mag above the single-star cooling track. Factoring in the uncertainty of a target’s Gaia HR diagram location from errors in the parallax and photometry, we slightly increase this minimum value and required the brightness of the binary to be 0.35 mag above the 0.6 WD cooling track to improve the search efficiency.
Going too far off of the cooling track would also contaminate our selection and its efficiency by introducing an abundance of WD–MS star candidates of the likes of cataclysmic variables (e.g. Abril et al., 2020), and with this in mind we limited the selection to systems with an absolute magnitude less than 1.2 mag brighter than the 0.6 WD cooling track. This upper boundary coincides with the typical cooling track for a helium-core WD of mass 0.3 , hence excluded DWD binaries from this upper brightness limit are those with an approximately 0.35 or less massive component. The cuts applied to the Gentile Fusillo et al. (2019) catalogue conclude as written in Table 1.
Starting DR2 WD catalogue: 486 641 sources | |
Selection for creating spline: | |
(Pwd > 0.3) and (Gmag < 18) and (Plx > 6.*e_Plx) | |
22 107 sources | |
Sample 20 evenly spaced bins between BP_RP = 0.65 & | |
1.65, fit with a spline spl_MG = f(BP_RP). | |
Select for a ‘clean’ sample of candidate targets: | |
(Gmag < 18.) and (Plx > 10.*e_Plx) | (1) |
22756 sources | |
(M_G < spl_MG0.35) and (M_G > spl_MG1.2) | (2) |
3494 sources | |
(BP_RP > 0.3) | (3) |
2697 sources | |
Select for an observable sample on the WHT telescope: | |
(Gmag < 17.) | (4) |
625 sources | |
() | (5) |
399 sources |
After applying a selection cut of Gmag < 17 to select targets that would grant a signal-to-noise (S/N) ratio greater than 15 per spectral element at the centre of H in less than a 30 min exposure time with the 4.2 m William Herschel Telescope (WHT, see Section 3), we were left with 625 sources. Lastly, we selected targets that were visible on site for prolonged durations ( > 20∘) to arrive at a final sample that was suitable for identification spectra of 399 sources.
We note that after applying the same cuts to Gaia DR3, each step of the sample selection had the following number of sources utilising the same numbering scheme: 1) 26 616, 2) 3729, 3) 2862, 4) 695, 5) 451. The primary cause of the increase in size of the final sample stems from the improved parallax uncertainties in Gaia DR3.
2.2 Sample biases
The sample cuts outlined in steps (1) to (3) in Table 1 cause the sample to be slightly biased towards brighter systems owing to the fact that brighter systems have more precise astrometric solutions, hence smaller parallax uncertainties. This is particularly relevant given the initial Gaia DR2 selection of the candidate sample, but far less so for Gaia DR3. The fact that the selection relies on a 0.6 M⊙ WD cooling track and that this is the median WD mass induces a natural bias towards at least one of the components being a WD of mass 0.6 M⊙ (). This is primarily the case for wide binaries where the two stars evolve in near-isolated conditions. WDs with smaller masses have larger radii and are hence more luminous for all effective temperatures compared to a WD with higher mass. With this and the fact that the upper MG selection cut of 1.2 mag above the 0.6 M⊙ cooling track coincides with that of a 0.3 M⊙ helium-core WD, we expect no DWD binaries that host an extremely-low-mass WD.
By applying a magnitude limit to the sample in steps (1) and (4), we induce a strong bias towards the frequency of hotter systems in the sample, which is clearly noticeable by the grouping of candidate systems with mag in Fig. 1. The magnitude limit imposed in the spline creation barely induces bias in the sample due to the large number of sources used in the spline fit, and the premise of introducing the magnitude limit in the step of creating the sample was to ensure good photometry across the BP-, RP- and G-bands.
The target selection, especially near the lining of the Milky Way and towards the Galactic bulge, is minorly biased towards brighter systems due to the impact of reddening since we applied no reddening correction to the sample selection. The most distant sources in the sample are found at 300 pc and the mean distance is 120 pc (Fig. 2), meaning that the impact of extinction is small and that even though the targets in our selection cuts will shift in colour to become bluer where some may lie slightly outside of the selection, the number that will is very small.
Limiting to a northern declination in step (5) is suspected to have negligible effects owing to the similarity between the results of volume-limited WD surveys in each hemisphere (McCleery et al., 2020; O’Brien et al., 2023). A final bias in the results of the targets in this sample, which does not stem from the sample selection itself, are introduced by the detection efficiency in the observing strategy (see Section 3).
3 Observations
With the final candidate sample outlined in Section 2.1, we obtained identification spectra of 117 candidates; 27 were observed for a single epoch, 9 for two epochs, and 88 candidates for 3 or more. Typically, more than 3 identification epochs were taken if the candidate was immediately identified as a promising double-lined candidate from the first 3 spectra. Our general observing strategy was to ensure that each candidate was observed on at least two separate nights spaced two nights apart; one of these nights with a single spectrum and the other night with two consecutive exposures to be sensitive to any ultra-compact systems with rapid RV variation. Generally speaking, compact DWDs have a median orbital period of approximately 1 day, meaning that the two day spacing granted sensitivity orbital periods up to a few days (Fig. 3), comprising over 90% of compact DWD systems (Nelemans et al., 2001).
These identification spectra were taken with the Intermediate-dispersion Spectrograph and Imaging System (ISIS) mounted on the 4.2 m William Herschel Telescope (WHT) across 20 nights, being 28 August 2018 - 5 September 2018, 13-15 February 2019, 19 February 2019, 15-17 April 2019, 10 & 11 June 2019 and 6 & 7 July 2019. ISIS utilises a dichroic filter to be able to observe in a blue and a red arm simultaneously, where we used the R600B grism for the blue arm centred on 4400 Å giving a wavelength coverage of 3770–5030 Å and the R1200R grism for the red arm centred on 6562 Å giving a wavelength coverage of 6250–6870 Å. We used a 1.0″slit width on all nights in 2018 and 1.1″for the others, having a spectral resolution of 0.81 Å and 0.74 Å in the red arm and 2.2 Å and 2 Å in the blue arm, respectively. A CuNe+CuAr arc exposure was taken for each new telescope pointing after acquiring a target for wavelength calibration, where the error in wavelength calibration for each arc frame in the blue arm was approximately 3 km s-1 and 2 km s-1 in the red arm. The science exposure time was always maintained at under 30 min not to suffer greatly from orbital smearing and a S/N ratio greater than 25 in the wings and 15 in the line core of H was sought after.
Bias frames and dome flats were taken at the beginning of each night and used in the data reduction. Spectrophotometric standard stars were also taken at the start or the end of each night to flux calibrate the science exposures and correct for the instrumental response function. When multiple arc frames were taken for a target, the arc closest in time to the science frame was used for wavelength calibration. All data were reduced using the Molly package (Marsh, 2019) and spectra were extracted using the method outlined in Marsh (1989).
The detection probability of a candidate being RV variable or double-lined, based off of the identification spectra alone, is displayed in Fig. 3. We determine a single-lined system to be RV variable if there is significant deviation from the mean velocity of all exposures (Section 5.2.2), while a double-lined system with two stars that contribute a similar flux is identifiable if the difference in velocity of the two stars is above 60 km s-1 for the resolution at H.
4 WD-BASS
[…]
WD-BASS is divided into 4 primary branches: 1) one-star spectral line fitting, 2) two-star spectral line fitting, 3) simultaneous or independent photometric SED fitting, and 4) RV fitting with a fixed atmospheric solution. Each branch will now be explained.
Access to the code is public and a full description of the code’s usage in its most recent version can be found on github111https://github.com/JamesMunday98/WD-BASS.
4.1 One-star fitting
Single star fitting in WD-BASS is performed by normalising and fitting specific spectral lines of interest. The user specifies the sets of wavelength ranges on each edge of the line to be used for a linear normalisation and the wavelength range that is fitted. The resolution of the observations at the spectral line of choice is also user supplied. Multiple spectra can be fit simultaneously or independently. One has the capability to fix or vary an atmospheric solution, fix or vary the RV of a star for each spectrum or, if the system is in a binary and the user supplies an orbital period (P) and reference epoch (T0), one can solve for a Keplerian orbit by varying the semi-amplitude (K) and/or the systemic velocity offset (Vγ). In the latter case, the eccentricity is assumed to be zero with a circularised orbit. Multiple spectral lines in a single spectrum can be fit independently or simultaneously with a fixed/varied RV for each line.
A Markov Chain Monte Carlo (MCMC) algorithm is employed in the code using the emcee python package (Foreman-Mackey et al., 2013) to probe a user-supplied parameter space and maximise the likelihood (which is equal to ) between the set of normalised observations and the normalised model fits sequentially. The interpolated spectrum for a trial solution is convolved to match the resolution of the observations with a Gaussian kernel.
4.2 Two-star fitting
The binary star fitting module of WD-BASS functions exactly the same as single star fitting but with the combination of two different model stars that can be of any spectral type, each star having a unique RV. Similar to before, any combination of the semi-amplitude of each star, , or their velocity offset, , can be input as fixed/fitted constraints (the velocity offset is not the same as the systemic velocity and is unique for the each star because of gravitational redshift).
Two stars in a binary have unique atmospheric parameters and radii, therefore in a normalised spectrum there is a relative flux scaling between the two. WD-BASS includes the options to fit for the relative flux scaling component, for it to be fixed, or for the scaling to be inferred through the radius of the two stars with T--radius relationships. In its current version, the radius of a DA star is inferred using the evolutionary track models of Istrate et al. (2016) when , Althaus et al. (2013) when and the hydrogen-rich evolutionary sequences of Bédard et al. (2020) otherwise, the assumption being that a star is likely a He-core WD and larger masses have a carbon-oxygen core. Hybrid carbon-oxygen/helium core composition WDs may exist in the mass range of – M⊙ (Zenati et al., 2019), or the hydrogen envelope could be large in some cases (Romero et al., 2019), but a spectrum that is poorly fit in this mass range can be overcome by allowing the scaling factor to vary freely. That said, the scaling is always interpolated in this study as these issues never arose in the fitting, which has often been the case for detached systems (Parsons et al., 2017). For helium-atmosphere DB/DC WDs, WD-BASS currently relies on a linear interpolation of the He-rich evolutionary sequences of Bédard et al. (2020) to predict the radius of the WD (M).
The model Eddington flux is converted to a flux observed at Earth () with the stellar radius (, inferred from –T–radius relationships) and the parallax, following
where , is the wavelength, is the flux and subscripts 1, 2 represent the brighter and dimmer star, respectively. The fluxes themselves are dependent on T, , T, .
4.3 Photometric SED fitting
In the advent of all-sky photometric surveys in the last decades, the astronomical community has a plethora of photometric magnitudes available in multiple wavebands that is publicly available and flux-calibrated. Yet, only recently have the parallaxes measured in the Gaia mission proved revolutionary in placing said stars distance scale (Gaia Collaboration et al., 2021, 2023). For sources up to mag in Gaia DR3 belonging to the 359 000 high-confidence WDs in the catalogue of Gentile Fusillo et al. (2021), 99.96% of the WDs have a parallax/parallax_error greater than 10, such that there is now the opportunity to exploit simultaneous fitting of spectral lines with well-flux-calibrated photometry to grant more precise atmospheric parameters. This is highly advantageous not only in single-lined systems, but especially so for double-lined binaries where the unique signature of each star may be difficult to disentangle, also allowing a more accurate system solution (Bédard et al., 2017). When handling photometry from satellite sources, air wavelengths in the model spectrum are converted to vacuum wavelengths.
To fit the photometry, synthetic spectra are created with equation 1 on an Eddington flux scale and can be reddened in WD-BASS with a supplied E(B-V) extinction constant and an assumed extinction curve with =3.1, following the extinction curves of Gordon et al. (2023). If fitting photometry simultaneously with spectroscopy, the same model is used for both datasets. The model is then integrated over the transmission spectrum of the respective filter to obtain the predicted observed flux in the bandpass. In the MCMC algorithm itself, a Gaussian prior on the parallax is applied and weighted by the supplied parallax error. The parallax, together with the radius inferred from and T, fixes the normalisation.
4.4 Improved radial velocity fitting
It is possible to determine RVs as discussed in the previous sub-sections and this is the standard procedure in WD-BASS. In case of a poor fit to the line core and specifically made for medium- to high-resolution data that includes H, where synthetic model spectra may struggle to reproduce the observed signature (Napiwotzki et al., 2020; Kilic et al., 2021), WD-BASS also includes the option to take an atmospheric solution and fit an extra Gaussian component to the star(s) model spectrum (spectra). This process is performed by taking the same set of spectra that were used in the one-star (two-star) atmospheric fitting, fixing the temperature(s), (s) and RVs found in the previously determined solution and then minimising the with an additional Gaussian solution for the star(s), having amplitude and standard deviation . An upper limit of Å is enforced to solely improve the line core, rather than any correction to the shape of the wings of the Balmer lines. Although there is no improvement to the atmospheric fit in the optionally employable method, the purpose of this step is to grant an improved RV accuracy and precision when relevant.
5 Methods: Fitting the candidate sample
5.1 Atmospheric parameters
For the fitting of all WD atmospheric classifications, we performed a bilinear interpolation in the T – space in trial solutions for DA WDs or trilinear in the T – –H/He space for DB WDs.
The observed spectra were normalised with a linear fit between the two wings of each spectral line and so were the model spectra (the spectral range for normalisation is different depending on the spectral type, detailed in Sections 5.1.1, 5.1.2). Our method in choosing the spectral region for normalisation was proportional to the temperature of the WD, with cooler WDs having less-broadened absorption lines. All data within the region of normalisation was included in a minimisation with a 4 outlier rejection threshold. We probed trial solutions using the MCMC algorithm with WD-BASS, utilising 100 walkers, a burn-in and a post-burn-in phase. In the case of double-lined systems where multiple exposures of a target were taken, we omit data where both stars are not visible in the spectrum in the atmospheric fitting (originating from an unfavourable orbital phase). In single-lined sources, we model all WHT identification spectra.
Multiple studies have shown a systematic difference between atmospheric constraints predominantly from fitting single WDs photometrically and spectroscopically that varies from survey to survey (Tremblay et al., 2019; Genest-Beaulieu & Bergeron, 2019; Gentile Fusillo et al., 2021; Cukanovaite et al., 2021; Sahu et al., 2023; Izquierdo et al., 2023), stemming from a difference in flux calibration between the surveys and imperfect model atmospheres. However, even with the systematics, all targets within our selection criteria are over-luminous compared to the 0.6 M⊙ WD cooling sequence and, in the vast majority of cases, the spectral features of one star dominate in the spectrum. We consider that the inclusion of photometry to reduce degeneracy in two-star atmospheric solutions outweighs the negatives induced from systematic effects in all cases. Therefore, we incorporate a hybrid fitting technique in WD-BASS of simultaneous spectroscopic and photometric fitting in all systems when available, whereby the combined with no weighting for the two datasets is minimised to reach a final solution.
We chose to limit the sources of photometry to Pan-STARRS DR1 (grizy) (Chambers & Pan-STARRS Team, 2018) and SDSS DR16 (ugriz) (Ahumada et al., 2020) and we chose not to include Gaia photometry as the very broad filters of Gaia are not optimal for two star fitting, where any extra flux calibration offset between Gaia and the other surveys could hamper the validity of the atmospheric solution. In any case, Gaia was still the best starting point for candidate selection due to its increased photometric precision and all-sky coverage. The extinction coefficient, AV, of each source was obtained through the reddening maps of Lallement et al. (2022) at a distance determined from the inverse of the Gaia DR3 parallax and was converted to before reddening the synthetic spectrum. An example showcasing the quality of fit to H–H8 following all of these steps with simultaneous photometric and spectroscopic hybrid fitting is displayed in Fig. 4 for one double-lined DWD.
5.1.1 DA WDs
Over the course of this study, we fitted the full sample of DWD binary candidates with two primary grids of model spectra: the 3D local thermodynamic equilibrium (LTE) DA spectra of Tremblay et al. (2013b, 2015) and the 1D non-LTE (NLTE) DA spectra of Kilic et al. (2021), which we will refer to as T15 and K21, respectively. Both grids rely on the line profiles of Tremblay & Bergeron (2009). The NLTE grid of K21 was computed by performing NLTE spectral synthesis on the LTE atmospheric structures of Tremblay & Bergeron (2009) using a modified version of the SYNSPEC code (see Kilic et al. 2021 for details). NLTE effects are especially important in the atmosphere of hot WDs (T 40 000 K; Napiwotzki 1997) but also remain non-negligible in the upper atmosphere of cooler objects, thus affecting the line cores (Heber et al., 1997; Koester et al., 1998). In particular, they are noticeable in high-resolution H observations of cool WDs in double-lined binaries (Bergeron et al., 1989; Kilic et al., 2021). The same effect of a decreased model flux in the line cores is apparent for the 3D vs 1D and NLTE vs LTE grids; in the former case, because of the cooling effects that come with the inclusion of convective overshoot (Tremblay et al., 2013b) and in the latter because of an increased opacity (increased NLTE lower level population) in the upper atmosphere of a WD where the lines are formed.
We observe that NLTE effects on H line cores are noticeable at medium resolution for the full range of effective temperatures and surface gravities in our sample. Both the grids of K21 and T15 performed well for all other Balmer lines observed at lower resolution, where the line core impact diminishes due to the increased broadening from the response of the blue grating. That said and even with the use of NLTE physics in the models, the synthetic line cores were too shallow. While the poorer fit caused in this case may not be detrimental to or hardly noticeable in single star fitting of an isolated WD, the impact is amplified in spectra with two DA WD components as the stars are blended, such that the line cores are the only components where a unique signature of each star is prominent and hence significantly impact the validity of the fit.
We were therefore encouraged to create a new synthetic spectral grid that approximately include both the 3D and NLTE effects on the line cores; a hybrid “3D-NLTE” grid that we introduce for the first time in this work. We started by dividing the NLTE and LTE grids of K21, which include otherwise identical input physics, yielding a NLTE correction factor . This factor was then interpolated onto the wavelength grid of T15 and applied to their 3D synthetic spectra. The final step was to broaden the product222Convective (turbulent) velocity broadening is not included by default in the synthetic spectral grid in T15. with the velocity vector in figure 11 of Tremblay et al. (2013a) at optical depth . This generated a NLTE correction to a 3D grid with resulting synthetic spectra (that will herein be referred to as 3D-NLTE) that incorporate 3D and NLTE physics without a full computational effort to include both effects simultaneously.
A comparison of T15 with the 3D-NLTE grid including simultaneous photometric fitting for three double-lined systems in the sample is displayed in Fig. 5. In addition, we compared the T15, K21 and 3D-NLTE grids by fitting our full observed sample with each. We found that the difference in the reported atmospheric parameters of each double-lined system is small ( K, dex for both stars) and smaller still for single-star fits to single-lined systems ( K, dex), yet we found that the 3D-NLTE grid outperformed the others by better fitting the observed spectra, especially at H. For that reason, we have utilised the 3D-NLTE spectral grid to fit all sources presented in this study. We note that for a minority of double-lined systems and specifically in the temperature range 9 000–10 000 K and surface gravity range 7.6–8.0 dex, we witness that the 3D-NLTE synthetic spectra slightly under-predict the flux at H compared to the observations; the synthetic line cores become too deep instead of too shallow (see e.g. WDJ212935.23+001332.26 in Appendix 8). We accredit the flux under-prediction to either small imperfections in the T15 grid since 3D effects are strongest in this temperature regime, or to an artefact of not creating a grid with the inclusion of 3D and NLTE effects being modelled together instead of our multiplicative factor approach.
In the process of atmospheric fitting and in general for DAs, we chose to linearly normalise Balmer lines using data in the range of km s-1 around H, H, H, H and km s-1 around H and H8 when T K. In WDs cooler than 8 000 K, inferred from the temperature quoted in Gentile Fusillo et al. (2021) from the fitting of Gaia photometry, we reduced the normalisation range to be approximately km s-1 and km s-1, for the respective groups of Balmer lines. The narrower prescription for cooler WDs was chosen to restrict the contribution of the continuum to the fit. Fine tuning of a couple of angstroms either side of the absorption features was performed to improve the normalisation and spectroscopic fit if a flux-calibration artefact was within the normalisation range.
Two DA systems in the sample are unique from the rest in the way that their spectra show a sharp emission at H only originating from background emission in the galactic plane, not from the system itself. These are WDJ183442.33+170028.00 and WDJ193845.80+264751.85, and to perform atmospheric fitting we mask the narrow hydrogen emission in both sources (see WDJ183442.33+170028.00 in Fig. 8). No photometric data is fitted for these sources.
5.1.2 Non-DA WDs
For the couple of DB WDs in our observed sample, we utilised synthetic spectra from Cukanovaite et al. (2021), incorporating 3D effects when possible. The observed and synthetic spectra were continuum normalised and all visible helium absorption features fitted. When including a helium-rich atmosphere DC WD component, we use the same synthetic spectra from Cukanovaite et al. (2021) but with a fixed H/He fraction of 10-5 (Bergeron et al., 2019; McCleery et al., 2020) and normalise spectra with a DC in the same way as for DA WDs. Hydrogen-rich atmosphere DC WDs exist below approximately 5 000 K and we continue using the T15 synthetic spectra to model hydrogen-rich DCs.
Some systems within our survey are not DA/DB/DC WDs (see Table 7). For the sample of interest, these systems are viewed as contaminants and as such we do not provide a full analysis of their atmospheric constraints in this paper. Notable targets of interest were 1) WDJ002215.19+423642.15 having a faint and shifted H towards the blue, perhaps indicative of a cool and magnetic star, 2) WDJ010343.47+555941.53 which could be an evolved CV or a subdwarf with an F/G/K star companion given the CH absorption band (G-band) at 4305Å (in Gaia DR3, this source has no parallax or proper motion, whereas it does in Gaia DR2). An analysis of time-series data is particularly encouraged for the last two systems.
5.1.3 Fitting single-lined DWD candidates
Irrelevant of the spectral type, our full observed sample is composed of DWD candidates. When we see unique signatures of both stars in the spectra and can identify individual spectral types, assigning the relevant atmosphere is simple. On the other hand, single-lined targets may still have an additional flux component that jeopardises the accuracy of a single-star solution.
To combat this dilemma, we perform two-star fits to all single-lined DA targets that show a missing flux component of a second WD, noticeable from overly shallow model fluxes in the line cores of the Balmer lines, a poor fit to the photometry or a fitted parallax outlier that deviates more than 3 from the expected value. The vast majority of targets from the DBL survey show Balmer lines, and it is expected that in a poorly-fitting target for a single-star DA model that the missing component is another DA WD. Hence, we assume that the most common companion would be another star of type DA. The second option that we consider is that the companion is a DC WD with a helium-rich atmosphere333A hydrogen atmosphere DC companion is considered when fitting DA synthetic spectra when the temperature is less than approximately 5 000 K as DAs become DCs and a H/He fraction (by number) of 10-5 (Cukanovaite et al., 2021). It is expected that a DC WD would have dex and we fix the surface gravity of the DC to remove parameter degeneracy in the fitting, before comparing a DA+DA and DA+DC () best-fit by eye. A visual comparison then allowed us to exclude a combination, if possible. As a last and separate check, we allowed the surface gravity of the DC to be free and comment on whether a DC WD with 7.5–9.0 would reproduce the observations and note if this is the case.
For any target where a single-star model fits well and we see little to no improvement in the of the hybrid fitting when including a second star, we consider the situation entirely degenerate and report an atmospheric solution with just the one WD, but we advise that caution is taken when interpreting the solution. If these systems are binaries, the temperature of the brighter component will be slightly overestimated and the surface gravity slightly underestimated.
5.2 Radial velocities
5.2.1 Fitting
To determine RVs, we fitted to H alone because of the much higher resolution in this spectral range for our observations, whereas the addition of data from the lower-resolution blue arm data would add confusion noise. Occasionally, and especially when the observing conditions worsened, the continuum S/N ratio can be relatively low (S/N<20), leading to a poor spectrum normalisation. We use the best atmospheric fit rather than the data to calculate the normalisation factors for the observed spectrum and then minimise the in the wavelength range within Å of H. Applying this re-normalisation slightly improved the accuracy and precision of the RVs in general and noticeably for the low S/N observations. Data within the Å range was sigma-clipped with a threshold of 4; excluding data near the line cores themselves. RVs and RV errors are reported from the median and standard deviation of 1000 bootstrapping iterations.
In double-lined DWD and single WD fitting, the model flux from each star is used in search for independent RVs. For the RVs of single-lined DWDs, a wide Gaussian component (to cover the broad wings of H) and a narrow Lorentizian component (to model the NLTE core) with a common RV were fit to all spectra to get a best model, and this model was then maintained fixed to search for independent RVs.
5.2.2 RV variability
From the identification spectra of double-lined systems, we can use the atmospheric solution and the measured RV of each component to assign each system a maximum orbital period. If we assume the unlikely case that the epoch of observation for all double-lined detection occurred when the stars were moving at a maximum orbital velocity and that the binary is angled edge-on (), the binary mass function can be rearranged to be
where we can take a rudimentary assumption that and are half of the maximum RV difference of the two stars. With an atmospheric solution bringing photometric masses of M1 and M2, the maximum period can be solved for.
While spotting a double-lined signature is often obvious upon inspection of the data, a statistically significant RV variability for a single-lined system is not. To search for single-lined RV variability the mean of all extracted RVs for a source is taken and a null-hypothesis that the RV is a constant with respect to the mean is tested. We compute the of all measurements compared to the mean and use the relevant -distribution for the number of degrees of freedom to calculate the probability that a source is not RV variable. For a system to be considered a candidate RV variable system, we require that it passes the 1% threshold and 0.01% to be considered as a WD binary with an unseen companion. All double-lined systems are naturally RV variable. All RV errors were propagated with an additional wavelength calibration error of 2 km s-1 in quadrature (Section 3).
6 Results
WDJ name | T | g1 | M1 | T | g2 | M2 | MT | D | Exp | RV | P | Ref |
[kK] | [dex] | [] | [kK] | [dex] | [] | [] | [pc] | # | [km s-1] | [d] | ||
J000319.54+022623.28 | 18.3 | 7.73 | 0.48 | 8.0 | 7.53 | 0.40 | 0.88 | 158.3 | 3 | 3.1 | - | |
J002602.29103751.86 | 9.9 | 7.55 | 0.41 | 5.4 | 8.12 | 0.66 | 1.07 | 88.5 | 3 | 0.6 | - | |
J005413.14+415613.73 | 7.7 | 7.77 | 0.47 | 7.4 | 7.65 | 0.44 | 0.90 | 54.1 | 2 | 1.4 | - | |
J013446.42+282616.83 | 13.7 | 7.77 | 0.49 | 9.7 | 7.60 | 0.43 | 0.92 | 177.2 | 3 | 1.0 | - | |
J013812.93+444252.10 | 15.0 | 7.92 | 0.57 | 8.1 | 7.88 | 0.53 | 1.10 | 81.6 | 3 | 2.5 | - | |
J014202.72+262354.58 | 12.3 | 7.87 | 0.54 | 8.4 | 7.68 | 0.43 | 0.96 | 173.2 | 4 | 1.4 | - | |
J020119.40050748.59 | 8.4 | 7.81 | 0.49 | 6.5 | 7.88 | 0.52 | 1.01 | 85.1 | 3 | 1.7 | - | |
J020847.22+251409.97 | 21.2 | 8.03 | 0.65 | 11.6 | 7.76 | 0.48 | 1.12 | 39.1 | 4 | 3.1 | - | |
J080856.79+461300.08 | 14.0 | 8.01 | 0.62 | 10.0 | 7.72 | 0.45 | 1.07 | 118.2 | 3 | 4.3 | - | |
J084457.81+453632.94 | 9.3 | 7.76 | 0.47 | 5.6 | 8.11 | 0.65 | 1.12 | 60.7 | 4 | 2.4 | - | |
J114446.16+364151.13 | 14.3 | 7.72 | 0.47 | 12.1 | 7.60 | 0.41 | 0.88 | 89.7 | 3 | 1.7 | - | |
J130014.82+181734.41 | 10.3 | 8.03 | 0.62 | 7.4 | 8.00 | 0.59 | 1.21 | 84.3 | 3 | 2.3 | - | |
J135342.35+165651.75 | 9.3 | 7.64 | 0.41 | 7.8 | 7.81 | 0.49 | 0.90 | 102.2 | 3 | 2.0 | - | |
J141625.94+311600.55 | 13.8 | 7.86 | 0.53 | 12.4 | 7.57 | 0.43 | 0.96 | 115.7 | 5 | 1.9 | - | |
J141632.84+111003.85 | 10.2 | 7.64 | 0.42 | 7.5 | 7.70 | 0.43 | 0.85 | 129.3 | 3 | 2.0 | - | |
J151109.90+404801.18 | 8.7 | 7.80 | 0.49 | 7.7 | 8.02 | 0.61 | 1.10 | 55.0 | 2 | 1.6 | - | |
J152038.37+390349.32 | 8.7 | 7.64 | 0.44 | 5.6 | 8.03 | 0.60 | 1.04 | 94.4 | 3 | 0.7 | - | |
J153615.83+501350.98 | 9.6 | 7.79 | 0.48 | 7.4 | 7.72 | 0.44 | 0.93 | 68.2 | 2 | 4.0 | 1 | |
J160822.19+420543.44 | 14.0 | 7.93 | 0.57 | 11.0 | 7.71 | 0.45 | 1.03 | 43.3 | 5 | 3.1 | 2 | |
J163441.85+173634.09 | 11.5 | 7.78 | 0.49 | 8.1 | 7.55 | 0.40 | 0.89 | 25.6 | 3 | 2.7 | 3 | |
J165935.59+620934.03 | 13.8 | 8.17 | 0.71 | 10.3 | 7.90 | 0.55 | 1.26 | 111.8 | 3 | 6.0 | - | |
J170120.99191527.57 | 19.2 | 7.74 | 0.49 | 13.9 | 8.33 | 0.82 | 1.31 | 97.0 | 3 | 0.4 | - | |
J180115.37+721848.76 | 18.1 | 7.96 | 0.60 | 10.9 | 7.80 | 0.49 | 1.09 | 128.4 | 3 | 1.8 | - | |
J180150.89+103401.08 | 22.4 | 8.14 | 0.71 | 11.4 | 7.78 | 0.49 | 1.20 | 115.9 | 2 | 8.3 | - | |
J181058.67+311940.94 | 20.2 | 8.16 | 0.72 | 16.5 | 8.35 | 0.83 | 1.55 | 49.0 | 6 | 1.9 | 4* | |
J182606.04+482911.30 | 14.4 | 7.84 | 0.53 | 10.9 | 7.64 | 0.42 | 0.95 | 136.0 | 2 | 3.4 | - | |
J183442.33170028.00 | 8.2 | 7.67 | 0.42 | 6.9 | 7.51 | 0.38 | 0.81 | 96.7 | 6 | 0.4 | - | |
J192002.51184442.99 | 20.1 | 8.21 | 0.75 | 11.9 | 7.98 | 0.59 | 1.34 | 155.7 | 2 | 1.3 | - | |
J192420.74+070135.14 | 16.4 | 8.06 | 0.65 | 14.4 | 7.90 | 0.56 | 1.21 | 161.7 | 3 | 2.6 | - | |
J211327.98+720814.03 | 11.5 | 7.80 | 0.50 | 7.5 | 7.42 | 0.37 | 0.87 | 96.2 | 3 | 8.4 | - | |
J212935.23+001332.26 | 9.2 | 7.68 | 0.43 | 8.2 | 7.68 | 0.42 | 0.85 | 65.4 | 2 | 0.9 | - | |
J214323.95175413.00 | 14.5 | 8.19 | 0.73 | 13.6 | 7.89 | 0.55 | 1.28 | 119.2 | 3 | 6.4 | - | |
J221209.01+612906.96 | 8.1 | 7.87 | 0.52 | 6.9 | 7.96 | 0.57 | 1.09 | 64.5 | 3 | 13.5 | - | |
J231404.30+552814.11 | 13.2 | 7.89 | 0.55 | 8.1 | 7.57 | 0.41 | 0.96 | 105.3 | 3 | 5.4 | - |
The best-fit atmospheric solutions are reported in Tables 2–6 as the 50th percentile of the post-burnin chains and errors as the 16th and 84th percentiles of this distribution. Within, subscripts 1 and 2 represent the hotter and cooler star, respectively.
34 systems are double-lined DWDs. Another 11 are double-lined DWD candidates as they show a small hint of a double-lined signature, however we can not confirm these as double-lined DWDs due to the S/N ratio of the data or perhaps an unfavourable orbital phase sampling. 27 single-lined DWDs were found where one star is visible in the spectrum but the source require an additional flux component for spectroscopic and photometric consistency, 38 sources appear as a single-lined DA WD with no companion present and 7 non-DA sources were found (Table 7). The location of these systems on the HR diagram separated by category is depicted in Fig. 6.
As the spectroscopic surface gravity of each star is largely defined by the relative depth of all Balmer lines compared to one another (Tremblay & Bergeron, 2009), we emphasise that care should be taken when interpreting the parameters of hybrid fits for stars with a temperature below approximately 5 500 K as the only Balmer lines visible in the data are H and at times H; all other Balmer lines vanish at lower temperatures. Regardless, the temperatures of the stars are expected to be accurate given the inclusion to the fit of flux-calibrated photometry. The error on these stars stays relatively low because the surface gravity is also constrained with the photometric data.
6.1 Double-lined DA DWDs
34 of our targets listed in Table 2 were found to be double-lined, double-DA DWDs, signifying at least a 29% detection efficiency in the survey. A further 3 targets found in Table 3 are deemed to be likely double-lined DWDs but the S/N ratio of the data leads to a lack of certainty, and a further 8 targets show a faint hint of a double-lined signature. If all are double-lined DWDs, the detection efficiency of the survey rises to 32% and 38%, respectively. These detection efficiencies are minima given the fact that the inclusion of Gaia DR3 parallaxes and colours have removed some systems from the original sample selection criteria. In the case that these other 11 candidates are not double-lined DWDs, they would be single-lined DWD binaries owing to the photometric and spectroscopic fit requiring the flux of a second DA WD for a good Balmer line and/or photometric fit, which then could be compact or unresolved, wide binaries.
Since our initial identification spectra, a couple of systems have since been discovered by other authors and shown to be double-lined DWDs. WDJ153615.83+501350.98 was earlier noted to be a double-degenerate candidate by Zuckerman et al. (2003), but since has been independently confirmed and discovered by Kilic et al. (2021, WD 1534+503) to be a double-lined DWD with much higher resolution observations at H. Using the WHT identification spectrum where the two stars are most separated (see Fig. 9), we obtain a 1 consistent atmospheric solution with the solution in Kilic et al. (2021) where our model reproduces H–H8, the photometric SED and the Gaia parallax excellently and H well. Kilic et al. (2021) do not fit to any data at H for atmospheric parameters and use H–H8, nevertheless, we recommend the usage of the system parameters described in their study owing to a slightly better consistency with the orbital solution for the photometric masses derived. Published in the same study, WDJ163441.85+173634.09 was independently discovered by Kilic et al. (2021, PG 1632+177) as a double-lined DWD and we again find 1-consistent atmospheric solution for both stars, where the fit in their study was also obtained using H–H8. WDJ160822.19+420543.44 was independently discovered by Kilic et al. (2020a, WD 1606+422) and we find a near-identical atmospheric solution for both stars as well. Besides these 3 double-lined systems, we believe that all of the other 31 double-lined DWDs presented in this study are new discoveries.
While not a clear double-lined DWD discovery, WDJ181058.67+311940.94 was flagged by Sahu et al. (2023) as a strong double-degenerate candidate from a poor spectral fit to its Lyman- profile. These authors do not extend their study to introduce two-star fits to the outliers in the data, meaning that the surface gravity of their single star solution is under-estimated and vastly different to our solution, but this is entirely expected due to the observed flux of the second star needing to be recovered by artificially inflating the radius in a single star fit. We have now confirmed the system as a DWD binary by detecting a double-lined feature at H.
Our solutions for the confirmed double-lined systems report that the masses of the hotter component range from 0.4–0.75 with a median mass of and that the dimmer companion has a median mass of . These measurements are slightly less than the 0.6 canonical mass of a WD, but this is expected given that most if not all of these compact double-lined DWDs have underwent a phase of mass loss in the past from binary interactions.
[…]
The new double-lined systems discovered seemingly show no preference for the cooler star being the more massive of the two. This challenges the scenario that the systems are formed via two common envelope phases that each led to significant shrinkage of the orbit, because in that case the more massive object would always have the largest cooling age. A feasible evolutionary path is instead that the less massive WD was first formed via stable mass transfer and cooled typically. Then, the other star evolved towards a WD, the system underwent a common envelope phase and exposed the star’s core. The new WD has a larger mass than the old one, but is still hotter being fresh out of the envelope, while the older, less-massive WD is cooler.
An interesting feature of the DBL survey is that we see a build-up of double-lined DWDs that runs approximately 0.5 mag above and parallel to the 0.6 M⊙ WD cooling track, as is prevalent in Fig. 6. The difference of 0.5 mag equates to roughly half of the flux, and so the natural conclusion is that double-lined DWD binaries are easiest to identify in our survey for two stars of similar brightness. While this is an unavoidable detection bias, this emphasises a strength of the DBL survey and an important point for the future of double-lined DWD detection outside of the DBL survey: 10s to 100s of double-lined DWDs in this strip of the HR diagram are undetectable from RV variability in lower resolution data where the two stars are inseparable and the two signals average to a net zero velocity change; at higher resolution, these are easy to identify.
6.2 Single-lined DWDs
27 sources in Table 5 were found to be single-lined, but a poor fit to the photometric and/or spectroscopic solution indicates that they are likely DWD binaries. Together with the double-lined systems and the double-lined candidate systems, this means that 72 of the 117 systems surveyed in the initial search selection using the Gaia DR2 astrometric solution data show a noticeable spectroscopic and/or photometric contribution from a WD companion.
Three of the single-lined DWDs, WDJ024323.67143957.37 (RV km s-1), WDJ172740.51+102337.94 (RV km s-1) and WDJ193833.62092519.87 (RV km s-1), show high RV variable log-probabilities of , and , respectively. Further spectroscopic observations of these three DWDs to trace the RV of the brighter star are encouraged to search for an orbital period. WDJ231519.82-052900.27 is a candidate for being RV variable with a log-probability of 2.3. All other sources with multiple spectra are consistent with a non-variable RV and most if not all are expected to be binaries with an orbital period greater than 10 days, where the only prospects of finding an orbital period are from a larger time baseline of astrometric or spectroscopic observations. Some of the non-RV variable candidates could have periods that fall in the aliases that we are not sensitive to in our search, which are predominantly aliases of 1 day (Fig. 3).
Overall in the sample of single- or double-lined DWDs, no system shows any signs of magnetism. If present, we would be sensitive to a magnetic field strength greater than approximately 50 kG with the identification spectra. The only DWD that has a strong magnetic field is NLTT 12758 (Kawka et al., 2017a), and the dynamo theory proposed by Schreiber et al. (2021) has been shown to support the lack of magnetic DWD binaries in the observed population and can explain the apparently chance timing of detecting the magnetic WD in NLTT 12758 (Schreiber et al., 2022). If true, a lack of magnetic DWDs in our survey is unsurprising, and the fact that we have specifically targeted systems with a 0.6 star and a dimmer, likely more massive, companion (much like the 0.83+0.69 pairing of NLTT 12758, Kawka et al., 2017a) alludes that we may have been more sensitive to a magnetic component with our survey criteria, but none were found.
6.3 Single star DA WDs
38 targets in Table 6 appear as single-lined, DA WDs where a single-star model fits the photometric and spectroscopic data well. In many of these systems, the reason is likely that they are isolated, low mass WDs or that one star strongly outshines its companion, which is highlighted by the grouping of single star DAs with mag in Fig. 6. Alternative reasons for identifying so many single star DA WDs in this region are that an additional unseen and red flux component makes the target appear redder and over-luminous, that the sources are indeed single WDs and a hybrid helium/carbon-oxygen core composition makes them appear over-luminous, that they are the product of an extremely-low-mass WD merger, or that some candidates are outliers in the Gaia database. Some of the sources classified as single DA WDs have crowded fields or issues with the Gaia DR3 parallax (see Table 6), meaning that they are spectroscopic fits instead of a hybrid. A few sources in Table 6 have M0.55 , and these are all cases where the location of the star on the HR diagram is on the lower edge of our search selection or when no photometry was fit. The minimum single star WD mass is approximately 0.3 , as was expected from the upper limit of the selection in Table 1. The median mass of the single star DA WDs in Table 6 is 0.43 .
We searched the single DA WDs for RV variations and a few targets displayed clear variability. The largest variation was in WDJ131913.72+450509.81 which exhibited an km s-1 shift between two exposures, strongly indicating that this source has an unseen component and is a WD binary. WDJ031242.85+221828.36 is also a clear case where an unseen component is present, having a maximum shift of km s-1. WDJ191927.69+395839.48 and WDJ230831.77+454212.19 exhibit a 30 km s-1 shift and both pass the threshold criteria of , making them RV variable systems. We promote follow-up observations of all to determine the orbital periods.
6.4 Gaia astrometric solutions
We cross-matched the Gaia non-single star catalogue (Halbwachs et al., 2023) with our full observed sample and we found that 5 in the observed sample have entries. One system that we flag as being a probable double-lined DWD but the S/N ratio leads to a lack of clarity, WDJ225123.02+293944.49, has an astrometric period of d, indicative that it may be a triple star system. The same can be said for the candidate double-lined DWD WDJ211345.93+262133.27 which has an astrometric period of 219.70.2 d. Three single-lined DWDs have detected astrometric periods, being WDJ023117.04+285939.88 ( d), WDJ205650.56+062149.68 ( d) and WDJ232519.87+140339.61 ( d). The most likely situation for these three is that they are wide DWDs and it is hence unsurprising that we do not witness a significant RV variability in any. Finally, WDJ232519.87+140339.61, a source that we have categorised as a single-lined WD with no significant flux contribution from a companion, has an astrometric period of d.
7 The fate of the observed DWD sample
[…]
The full series of photometric masses of double-lined DWD binaries identified in our work is plotted in Fig. 7, that shows the assessment of Shen (2015) of the most likely outcomes for mass combination but in many cases may be quite uncertain. Here and only in this figure, subscripts 1 and 2 are the more and less massive components, respectively. Single-lined DWDs are omitted since many could be wide binaries. As mentioned earlier, the vast majority of DWD binaries in the literature that constitute the observed population come in the form of low-mass WD binaries, which is far from reflective of the population as a whole (see e.g. Toonen et al., 2012). An immediate change in the observed population of DWDs is now obvious from our uncovered sample, with the majority of our systems having larger combined total masses. None of our sample falls into the low-mass regime and the majority of systems are dominated by a more massive star of approximately 0.55 .
[…]
WDJ name | T | g1 | M1 | T | g2 | M2 | MT | D | Exp | RV | P | Ref |
[kK] | [dex] | [] | [kK] | [dex] | [] | [] | [pc] | # | [km s-1] | [d] | ||
J062538.73162132.02 | 14.1 | 8.03 | 0.63 | 11.3 | 7.81 | 0.50 | 1.13 | 149.5 | 1 | 7.4 | - | |
J211927.07030002.38 | 9.4 | 7.70 | 0.44 | 5.4 | 8.52 | 0.92 | 1.36 | 93.7 | 3 | 0.6 | - | |
J225123.02+293944.49 | 5.7 | 7.96 | 0.56 | 5.6 | 8.52 | 0.92 | 1.49 | 20.0 | 3 | 5.4 | - |
WDJ name | T | g1 | M1 | T | g2 | M2 | MT | D | Exp | RV | Ref |
[kK] | [dex] | [] | [kK] | [dex] | [] | [] | [pc] | # | [km s-1] | ||
J053316.88075049.72 | 11.2 | 7.27 | 0.36 | 7.5 | 7.05 | 0.25 | 0.61 | 80.5 | 1 | - | |
J141354.17+153020.71 | 15.3 | 7.78 | 0.50 | 13.2 | 7.97 | 0.59 | 1.09 | 147.7 | 2 | - | |
J145011.93191408.67 | 8.2 | 7.85 | 0.51 | 6.2 | 7.70 | 0.43 | 0.94 | 48.6 | 5 | 1 | |
J180600.36002720.92 | 24.1 | 7.60 | 0.45 | 8.5 | 7.46 | 0.37 | 0.82 | 147.7 | 3 | - | |
J193642.54054744.38 | 10.6 | 7.99 | 0.59 | 7.9 | 7.83 | 0.50 | 1.09 | 105.3 | 3 | - | |
J204922.92+262517.50 | 16.9 | 7.83 | 0.53 | 14.6 | 8.06 | 0.65 | 1.17 | 141.7 | 4 | - | |
J211345.93+262133.27 | 8.7 | 7.76 | 0.47 | 5.6 | 7.88 | 0.52 | 0.99 | 31.6 | 3 | 2 | |
J234929.57+102255.57 | 14.5 | 7.80 | 0.50 | 7.0 | 7.83 | 0.50 | 1.00 | 117.3 | 3 | - |
WDJ name | T | g | T | g | T | g | T | D | Exp | p | Ref |
[kK] | [dex] | [kK] | [dex] | [kK] | [dex] | [kK] | [pc] | # | |||
J001353.60+204852.65 | 17.8 | 7.89 | 8.4 | 7.82 | 16.7 | 7.80 | 6.3 | 294.4 | 3 | -0.6 | - |
J003045.78+273333.36 | 16.2 | 7.90 | 7.9 | 7.74 | 14.7 | 7.74 | * | 164.0 | 4 | -0.2 | - |
J014511.23+313243.56 | 10.0 | 7.97 | 6.2 | 8.00 | 10.0 | 7.99 | 6.5 | 36.5 | 1 | - | 1 |
J023117.04+285939.88 | 7.5 | 8.01 | 5.8 | 7.99 | 8.0 | 8.13 | * | 27.9 | 5 | -0.0 | - |
J024323.67143957.37 | 9.0 | 7.63 | 5.0 | 7.86 | 9.8 | 7.92 | * | 79.4 | 3 | -99.0 | - |
J080739.31+132110.65 | 11.7 | 7.66 | 6.8 | 7.71 | 11.0 | 7.57 | * | 114.7 | 1 | - | - |
J084417.70+750008.79 | 14.6 | 8.07 | 6.9 | 8.00 | 13.4 | 7.95 | 6.3 | 101.7 | 4 | -0.0 | - |
J084634.41+194505.18 | 10.9 | 7.61 | 7.5 | 7.59 | 10.4 | 7.51 | * | 147.7 | 1 | - | - |
J102459.83+044610.50 | 12.3 | 7.61 | 5.7 | 7.66 | 12.4 | 7.61 | * | 43.2 | 1 | - | - |
J104709.19+345346.65 | 9.1 | 7.63 | 5.2 | 7.82 | 9.1 | 7.69 | * | 79.3 | 1 | - | - |
J113100.20+493826.27 | 10.9 | 7.75 | 6.9 | 8.03 | 10.8 | 7.71 | 6.5 | 98.7 | 1 | - | - |
J113347.81+624313.29 | 6.9 | 7.65 | 5.6 | 8.88 | 6.4 | 7.89 | 6.0 | 45.8 | 1 | - | 2 |
J131257.90+580511.29 | 11.6 | 8.01 | 6.4 | 8.05 | 11.4 | 8.00 | 6.4 | 30.9 | 2 | -0.0 | - |
J135738.69025819.41 | 16.9 | 7.75 | 14.1 | 8.23 | 15.7 | 7.54 | 6.6 | 136.9 | 2 | -0.2 | - |
J154214.21034131.29 | 11.6 | 8.16 | 8.1 | 7.83 | 11.0 | 7.95 | 5.3 | 53.1 | 1 | - | - |
J172740.51+102337.94 | - | - | - | - | 10.1 | 8.04 | 7.8 | 58.9 | 5 | -9.9 | - |
J191329.91+163822.12 | 8.4 | 7.94 | 6.7 | 7.91 | 8.9 | 8.03 | * | 77.0 | 5 | -1.0 | - |
J192359.24+214103.62 | 9.5 | 7.77 | 5.7 | 8.08 | 10.0 | 8.00 | 6.4 | 35.1 | 3 | -0.0 | - |
J193833.62092519.87 | 16.0 | 8.12 | 8.4 | 7.68 | 12.5 | 7.68 | 6.9 | 115.2 | 3 | -6.7 | - |
J204517.85+810503.40 | 8.6 | 8.16 | 7.6 | 7.97 | 9.3 | 8.12 | * | 67.0 | 3 | -1.7 | 1 |
J205650.56+062149.68 | 10.4 | 7.83 | 6.5 | 7.95 | 10.3 | 7.83 | 6.4 | 95.3 | 3 | -0.6 | - |
J213616.39131834.50 | 10.2 | 7.83 | 5.1 | 7.92 | - | - | - | 23.4 | 3 | -0.1 | - |
J214632.37+155039.08 | 8.6 | 7.89 | 6.8 | 7.96 | 9.1 | 7.93 | * | 72.8 | 3 | -0.2 | - |
J221052.87045540.80 | 7.9 | 7.86 | 6.8 | 7.87 | - | - | - | 63.4 | 3 | -0.1 | - |
J231443.05073417.85 | 14.1 | 8.10 | 9.6 | 7.93 | - | - | - | 75.2 | 3 | -0.1 | - |
J231519.82052900.27 | 6.6 | 7.69 | 5.7 | 7.71 | 6.5 | 7.63 | * | 54.7 | 3 | -2.3 | - |
J233041.67+110206.43 | 21.5 | 7.96 | 6.5 | 8.07 | 21.4 | 7.96 | 6.4 | 98.8 | 1 | - | - |
WDJ name | T | g | M | D | Spec | p | Ref | Other Name |
[kK] | [cm s-2] | [] | [pc] | # | ||||
J001321.07+282019.83 | 26.2 | 7.84 | 0.56 | 135.3 | 13 | -0.5 | 1 | PG 0010+281 |
J003508.27+135045.32 | 22.9 | 7.47 | 0.44 | 246.8 | 3 | -0.2 | 2 | WD 0032+135 |
J015437.45+374145.37 | 12.9 | 7.53 | 0.42 | 136.6 | 3 | -0.1 | - | - |
J022631.25+203106.04 | 21.5 | 7.89 | 0.57 | 164.9 | 3 | -0.1 | - | - |
J031242.85+221828.36 | 6.3 | 7.36 | 0.33 | 63.2 | 3 | -99.0 | 3, 4 | LP 355-39 |
J042127.88+701419.37 | 12.4 | 7.60 | 0.41 | 54.5 | 1 | - | 1 | GD 429 |
J074837.63+004011.72 | 15.5 | 7.57 | 0.40 | 184.9 | 3 | -0.8 | - | - |
J081706.44+054954.62 | 11.3 | 7.61 | 0.41 | 133.3 | 1 | - | - | - |
J091914.80+772350.79 | 9.1 | 7.56 | 0.41 | 46.5 | 6 | -0.0 | 3 | - |
J093709.98+650746.91 | 18.1 | 7.92 | 0.58 | 105.5 | 4 | -0.3 | - | SDSS J093709.96+650746.6 |
J101606.87011917.14 | 8.1 | 7.63 | 0.43 | 46.3 | 1 | - | 3, 5 | EGGR 2532 |
J114604.37+051401.54 | 6.6 | 7.43 | 0.35 | 62.3 | 1 | - | 3, 4, 5 | WD 1143+055 |
J130313.03032323.92 | 7.2 | 7.56 | 0.40 | 64.6 | 1 | - | 3, 5 | WD 1300-031 |
J131913.72+450509.81 | 13.1 | 7.43 | 0.40 | 49.0 | 2 | -99.0 | 5 | EGGR 188 |
J134503.00110434.10 | 21.0 | 7.57 | 0.42 | 246.6 | 2 | -1.0 | - | - |
J142047.04+465601.58 | 10.2 | 7.58 | 0.42 | 98.6 | 2 | -0.1 | 3, 5 | WD 1418+471 |
J150402.99+345440.76 | 18.8 | 8.07 | 0.66 | - | 6 | -0.2 | 1 | PG 1502+351 |
J152849.51012853.74 | 12.8 | 7.93 | 0.57 | 52.4 | 3 | -1.5 | - | - |
J155840.22+162556.04 | 12.3 | 7.47 | 0.40 | 106.3 | 1 | - | 1 | 2MASS J15584020+1625556 |
J160647.76152740.19 | 31.9 | 7.98 | 0.64 | 194.4 | 4 | -3.4 | - | - |
J175644.61020847.66 | 16.1 | 7.52 | 0.43 | 157.8 | 4 | -0.0 | - | - |
J181304.62+222449.79 | 21.4 | 7.91 | 0.58 | 209.3 | 2 | -0.8 | - | - |
J181339.31+255058.16 | 19.1 | 7.53 | 0.40 | 183.5 | 3 | -0.5 | - | - |
J182444.29+600159.40 | 12.7 | 7.70 | 0.45 | 96.2 | 3 | -0.1 | ||
J183752.49125257.45 | 37.8 | 8.21 | 0.78 | 179.5 | 3 | -0.2 | - | - |
J185640.86+120844.61 | 13.6 | 7.53 | 0.42 | 156.3 | 3 | -0.0 | - | - |
J190401.01102305.20 | 16.2 | 7.92 | 0.57 | 103.0 | 3 | -0.1 | - | - |
J191927.69+395839.48 | 18.9 | 7.73 | 0.49 | 105.8 | 3 | -15.1 | 6 | KIC 4829241 |
J192817.81+354442.60 | 20.5 | 7.92 | 0.58 | 138.2 | 1 | - | - | - |
J193845.80+264751.85 | 22.2 | 7.93 | 0.59 | 97.0 | 5 | -0.2 | - | - |
J195314.95105417.89 | 22.6 | 7.74 | 0.50 | 145.0 | 1 | - | - | - |
J195622.94+641359.19 | 18.3 | 7.77 | 0.50 | 85.5 | 3 | -0.2 | - | - |
J205020.65+263040.76 | 5.2 | 8.54 | 0.94 | 19.1 | 1 | - | 3, 4, 7, 8 | GJ 4166 |
J230831.77+454212.19 | 13.6 | 7.48 | 0.41 | 103.1 | 3 | -6.8 | - | - |
J231406.68+233343.06 | 7.1 | 7.45 | 0.36 | 68.5 | 3 | -0.2 | 3 | - |
J232519.87+140339.61 | 4.9 | 7.29 | 0.30 | 23.5 | 1 | - | 3, 4, 7, 8 | WD 2322+137 |
J232557.82+255222.39 | 5.6 | 7.49 | 0.38 | 46.0 | 3 | -0.0 | 3, 4 | EGGR 380 |
J235313.18+205117.58 | 7.4 | 7.64 | 0.43 | 57.0 | 3 | -0.2 | 3, 4 | WD 2350+205 |
WDJ name | D | SpT | Note |
J002215.19+423642.15 | 34.5 | DC | Perhaps H absorption at 6554Å. |
J004502.15040710.05 | 43.2 | DQ | Swan bands |
J010343.47+555941.53 | 26.1 | Not a WD | Evolved CV or sd+F/G/K star. Parallax in Gaia DR2 not DR3 |
J021243.27080216.23 | 119.7 | MS | No WD signatures. Common proper motion pair MS & blue object |
J182138.99+144158.19 | 89.7 | DB | T kK, dex |
J195513.90+222458.79 | 14.8 | MS | No signature of WD. Parallax in DR2 not DR3 |
J201437.22+231607.23 | 29.8 | MS | No signature of WD. Parallax in Gaia DR2 not DR3 |
8 Conclusions
We have presented a pilot study including the first results from the DBL survey based on 20 nights of observations with the WHT. We surveyed a large sample of 117 DWD binary candidates that reside above the 0.6 M⊙ WD cooling track through the exploitation of the Gaia DR2 HR diagram, randomly sampling candidates from a magnitude limited selection. Then, we fitted and obtained atmospheric solutions for the entire sample when all visible stars in the spectra were of spectral type DA, DB or DC with the custom-made fitting code WD-BASS that is designed for time-series spectroscopy of WDs and is publicly available for use.
The primary goal of the survey was to find double-lined DWDs which allow the obtention of precise masses for both stars in the binary through atmospheric fitting. Our methods demonstrate at least a 29% double-lined detection rate after a couple of observing epochs and that 73 of the 117 candidates show a separable spectroscopic/photometric flux contribution that must arise from a DWD configuration. A further 6 single-lined WDs are flagged for having an unseen companion based off RV variability. All double-lined DWDs are compact binaries, and the large RV difference between the two stars provides some insight to the orbital period distribution of the systems.
[…]
In continued work we will report orbital solutions of many of the double-lined DWDs in this study. Many more double-lined DWDs are waiting to be discovered within the search selection, where we will strive to obtain further completion of the 625 <17 mag double degenerate candidates.
Acknowledgements
[…]
Based on observations made with the William Herschel Telescope operated on the island of La Palma by the Isaac Newton Group of Telescopes in the Spanish Observatorio del Roque de los Muchachos of the Instituto de Astrofísica de Canarias.
Data Availability
The reduced spectra and extracted RVs are available upon request to the authors. Raw data can be obtained through the Isaac Newton Group data archive with the dates of observations mentioned in text. Our compilation of all discovered DWD binaries is available through Github at https://github.com/JamesMunday98/CloseDWDbinaries.
Appendix A All double-lined systems
All DWDs that exhibit a double-lined feature are displayed in Fig. 8 and all of the likely double-lined DWDs that should be confirmed with further observations are found in Figs. 9 and 10. While hybrid fitting was used to derive atmospheric parameters, which includes the simultaneous fitting of survey photometry and our spectra, the data and model spectrum at H alone are shown in these figures to emphasise the double-lined nature of each system.
Appendix B Sensitivity to double-lined systems
The ability to detect a double-lined DWD is dependent on the relative flux contribution of the two stars, and hence there is a dependence on the stars’ radii. We depict the impact of this effect in Fig. 12 to determine for which combinations of temperature and surface gravity would the companion contribute at least 25% of the flux at the centre of H in our data. A minimum flux contribution of 25% was used to construct the selection criteria in Section 2.1, and below this threshold a second star is unlikely to be detectable at the signal-to-noise ratio of our data.

### v3

@@
 [kK] | [dex] | [] | [kK] | [dex] | [] | [] | [pc] | # | [km s-1] | [d] | ||
-J000319.54+022623.28 | 18.3 | 7.73 | 0.48 | 8.0 | 7.53 | 0.40 | 0.88 | 158.3 | 3 | 3.1 | - | |
-J002602.29103751.86 | 9.9 | 7.55 | 0.41 | 5.4 | 8.12 | 0.66 | 1.07 | 88.5 | 3 | 0.6 | - | |
-J005413.14+415613.73 | 7.7 | 7.77 | 0.47 | 7.4 | 7.65 | 0.44 | 0.90 | 54.1 | 2 | 1.4 | - | |
-J013446.42+282616.83 | 13.7 | 7.77 | 0.49 | 9.7 | 7.60 | 0.43 | 0.92 | 177.2 | 3 | 1.0 | - | |
-J013812.93+444252.10 | 15.0 | 7.92 | 0.57 | 8.1 | 7.88 | 0.53 | 1.10 | 81.6 | 3 | 2.5 | - | |
-J014202.72+262354.58 | 12.3 | 7.87 | 0.54 | 8.4 | 7.68 | 0.43 | 0.96 | 173.2 | 4 | 1.4 | - | |
-J020119.40050748.59 | 8.4 | 7.81 | 0.49 | 6.5 | 7.88 | 0.52 | 1.01 | 85.1 | 3 | 1.7 | - | |
-J020847.22+251409.97 | 21.2 | 8.03 | 0.65 | 11.6 | 7.76 | 0.48 | 1.12 | 39.1 | 4 | 3.1 | - | |
-J080856.79+461300.08 | 14.0 | 8.01 | 0.62 | 10.0 | 7.72 | 0.45 | 1.07 | 118.2 | 3 | 4.3 | - | |
-J084457.81+453632.94 | 9.3 | 7.76 | 0.47 | 5.6 | 8.11 | 0.65 | 1.12 | 60.7 | 4 | 2.4 | - | |
-J114446.16+364151.13 | 14.3 | 7.72 | 0.47 | 12.1 | 7.60 | 0.41 | 0.88 | 89.7 | 3 | 1.7 | - | |
-J130014.82+181734.41 | 10.3 | 8.03 | 0.62 | 7.4 | 8.00 | 0.59 | 1.21 | 84.3 | 3 | 2.3 | - | |
-J135342.35+165651.75 | 9.3 | 7.64 | 0.41 | 7.8 | 7.81 | 0.49 | 0.90 | 102.2 | 3 | 2.0 | - | |
-J141625.94+311600.55 | 13.8 | 7.86 | 0.53 | 12.4 | 7.57 | 0.43 | 0.96 | 115.7 | 5 | 1.9 | - | |
-J141632.84+111003.85 | 10.2 | 7.64 | 0.42 | 7.5 | 7.70 | 0.43 | 0.85 | 129.3 | 3 | 2.0 | - | |
-J151109.90+404801.18 | 8.7 | 7.80 | 0.49 | 7.7 | 8.02 | 0.61 | 1.10 | 55.0 | 2 | 1.6 | - | |
-J152038.37+390349.32 | 8.7 | 7.64 | 0.44 | 5.6 | 8.03 | 0.60 | 1.04 | 94.4 | 3 | 0.7 | - | |
-J153615.83+501350.98 | 9.6 | 7.79 | 0.48 | 7.4 | 7.72 | 0.44 | 0.93 | 68.2 | 2 | 4.0 | 1 | |
-J160822.19+420543.44 | 14.0 | 7.93 | 0.57 | 11.0 | 7.71 | 0.45 | 1.03 | 43.3 | 5 | 3.1 | 2 | |
-J163441.85+173634.09 | 11.5 | 7.78 | 0.49 | 8.1 | 7.55 | 0.40 | 0.89 | 25.6 | 3 | 2.7 | 3 | |
-J165935.59+620934.03 | 13.8 | 8.17 | 0.71 | 10.3 | 7.90 | 0.55 | 1.26 | 111.8 | 3 | 6.0 | - | |
-J170120.99191527.57 | 19.2 | 7.74 | 0.49 | 13.9 | 8.33 | 0.82 | 1.31 | 97.0 | 3 | 0.4 | - | |
-J180115.37+721848.76 | 18.1 | 7.96 | 0.60 | 10.9 | 7.80 | 0.49 | 1.09 | 128.4 | 3 | 1.8 | - | |
-J180150.89+103401.08 | 22.4 | 8.14 | 0.71 | 11.4 | 7.78 | 0.49 | 1.20 | 115.9 | 2 | 8.3 | - | |
-J181058.67+311940.94 | 20.2 | 8.16 | 0.72 | 16.5 | 8.35 | 0.83 | 1.55 | 49.0 | 6 | 1.9 | 4* | |
-J182606.04+482911.30 | 14.4 | 7.84 | 0.53 | 10.9 | 7.64 | 0.42 | 0.95 | 136.0 | 2 | 3.4 | - | |
-J183442.33170028.00 | 8.2 | 7.67 | 0.42 | 6.9 | 7.51 | 0.38 | 0.81 | 96.7 | 6 | 0.4 | - | |
-J192002.51184442.99 | 20.1 | 8.21 | 0.75 | 11.9 | 7.98 | 0.59 | 1.34 | 155.7 | 2 | 1.3 | - | |
-J192420.74+070135.14 | 16.4 | 8.06 | 0.65 | 14.4 | 7.90 | 0.56 | 1.21 | 161.7 | 3 | 2.6 | - | |
-J211327.98+720814.03 | 11.5 | 7.80 | 0.50 | 7.5 | 7.42 | 0.37 | 0.87 | 96.2 | 3 | 8.4 | - | |
-J212935.23+001332.26 | 9.2 | 7.68 | 0.43 | 8.2 | 7.68 | 0.42 | 0.85 | 65.4 | 2 | 0.9 | - | |
-J214323.95175413.00 | 14.5 | 8.19 | 0.73 | 13.6 | 7.89 | 0.55 | 1.28 | 119.2 | 3 | 6.4 | - | |
-J221209.01+612906.96 | 8.1 | 7.87 | 0.52 | 6.9 | 7.96 | 0.57 | 1.09 | 64.5 | 3 | 13.5 | - | |
-J231404.30+552814.11 | 13.2 | 7.89 | 0.55 | 8.1 | 7.57 | 0.41 | 0.96 | 105.3 | 3 | 5.4 | - |
+J000319.54+022623.28 | 18.2 | 7.69 | 0.47 | 7.5 | 7.48 | 0.38 | 0.85 | 158.3 | 3 | 3.0 | - | |
+J002602.29103751.86 | 10.7 | 7.74 | 0.47 | 5.8 | 7.60 | 0.42 | 0.88 | 88.3 | 3 | 1.2 | - | |
+J005413.14+415613.73 | 7.7 | 7.69 | 0.43 | 7.4 | 7.73 | 0.45 | 0.88 | 54.0 | 2 | 1.2 | - | |
+J013446.42+282616.83 | 13.1 | 7.60 | 0.41 | 9.3 | 7.72 | 0.45 | 0.86 | 176.3 | 3 | 0.7 | - | |
+J013812.93+444252.10 | 15.0 | 7.92 | 0.57 | 8.0 | 7.88 | 0.53 | 1.10 | 81.5 | 3 | 2.4 | - | |
+J014202.72+262354.58 | 12.2 | 7.86 | 0.53 | 8.3 | 7.72 | 0.45 | 0.97 | 171.4 | 4 | 1.3 | - | |
+J020119.40050748.59 | 8.3 | 7.80 | 0.49 | 6.7 | 7.91 | 0.54 | 1.02 | 85.0 | 3 | 1.6 | - | |
+J020847.22+251409.97 | 21.2 | 7.86 | 0.55 | 11.6 | 8.21 | 0.74 | 1.29 | 39.1 | 4 | 1.5 | - | |
+J080856.79+461300.08 | 14.0 | 7.99 | 0.60 | 10.1 | 7.76 | 0.47 | 1.08 | 118.2 | 3 | 4.0 | - | |
+J084457.81+453632.94 | 9.8 | 7.97 | 0.58 | 5.9 | 7.71 | 0.43 | 1.01 | 60.9 | 4 | 5.7 | - | |
+J114446.16+364151.13 | 13.3 | 7.63 | 0.42 | 13.0 | 7.69 | 0.45 | 0.87 | 89.7 | 3 | 1.3 | - | |
+J130014.82+181734.41 | 10.9 | 8.13 | 0.68 | 6.7 | 7.75 | 0.46 | 1.14 | 84.3 | 3 | 3.5 | - | |
+J135342.35+165651.75 | 9.6 | 7.76 | 0.47 | 7.6 | 7.70 | 0.43 | 0.90 | 101.4 | 3 | 2.9 | - | |
+J141625.94+311600.55 | 13.3 | 7.74 | 0.47 | 12.8 | 7.62 | 0.42 | 0.89 | 115.5 | 5 | 1.6 | - | |
+J141632.84+111003.85 | 10.5 | 7.76 | 0.47 | 7.5 | 7.60 | 0.42 | 0.89 | 129.1 | 3 | 2.6 | - | |
+J151109.90+404801.18 | 9.1 | 8.12 | 0.67 | 7.6 | 7.71 | 0.44 | 1.11 | 54.9 | 2 | 4.1 | - | |
+J152038.37+390349.32 | 9.6 | 8.02 | 0.61 | 5.4 | 7.35 | 0.32 | 0.93 | 94.4 | 3 | 2.1 | - | |
+J153615.83+501350.98 | 9.6 | 7.73 | 0.46 | 7.3 | 7.79 | 0.48 | 0.94 | 68.2 | 2 | 3.3 | 1 | |
+J160822.19+420543.44 | 14.0 | 7.89 | 0.55 | 11.0 | 7.76 | 0.47 | 1.02 | 43.3 | 5 | 2.7 | 2 | |
+J163441.85+173634.09 | 11.4 | 7.69 | 0.44 | 8.1 | 7.80 | 0.48 | 0.93 | 25.6 | 3 | 1.9 | 3 | |
+J165935.59+620934.03 | 13.0 | 7.80 | 0.50 | 8.2 | 8.17 | 0.70 | 1.20 | 110.9 | 3 | 2.3 | - | |
+J170120.99191527.57 | 20.5 | 8.08 | 0.67 | 13.5 | 7.75 | 0.48 | 1.15 | 96.7 | 3 | 1.1 | - | |
+J180115.37+721848.76 | 18.0 | 7.88 | 0.55 | 11.2 | 8.02 | 0.62 | 1.17 | 128.0 | 3 | 1.3 | - | |
+J180150.89+103401.08 | 22.0 | 7.92 | 0.59 | 8.3 | 7.89 | 0.53 | 1.12 | 114.9 | 2 | 5.4 | - | |
+J181058.67+311940.94 | 20.0 | 8.16 | 0.72 | 16.7 | 8.35 | 0.83 | 1.55 | 48.9 | 6 | 1.9 | 4* | |
+J182606.04+482911.30 | 14.4 | 7.72 | 0.47 | 11.3 | 7.89 | 0.54 | 1.01 | 135.6 | 2 | 2.1 | - | |
+J183442.33170028.00 | 8.2 | 7.59 | 0.42 | 7.0 | 7.76 | 0.46 | 0.88 | 96.7 | 6 | 0.3 | - | |
+J192002.51184442.99 | 20.4 | 8.17 | 0.73 | 12.0 | 8.08 | 0.65 | 1.38 | 156.1 | 2 | 1.1 | - | |
+J192420.74+070135.14 | 16.7 | 7.96 | 0.59 | 14.1 | 8.03 | 0.63 | 1.22 | 161.0 | 3 | 1.9 | - | |
+J211327.98+720814.03 | 11.1 | 7.63 | 0.42 | 7.0 | 7.50 | 0.38 | 0.79 | 96.4 | 3 | 6.0 | - | |
+J212935.23+001332.26 | 9.2 | 7.69 | 0.44 | 7.9 | 7.64 | 0.44 | 0.87 | 65.5 | 2 | 0.9 | - | |
+J214323.95175413.00 | 14.2 | 8.05 | 0.64 | 14.0 | 8.04 | 0.64 | 1.28 | 118.2 | 3 | 4.3 | - | |
+J221209.01+612906.96 | 8.1 | 7.90 | 0.54 | 7.0 | 7.93 | 0.55 | 1.09 | 64.5 | 3 | 14.9 | - | |
+J231404.30+552814.11 | 14.0 | 8.08 | 0.66 | 8.6 | 7.50 | 0.38 | 1.04 | 105.6 | 3 | 7.9 | - |
 The best-fit atmospheric solutions are reported in Tables 2–6 as the 50th percentile of the post-burnin chains and errors as the 16th and 84th percentiles of this distribution.
@@
 These measurements are slightly less than the 0.6 canonical mass of a WD, but this is expected given that most if not all of these compact double-lined DWDs have underwent a phase of mass loss in the past from binary interactions.
-[…]
+WDJ183442.33170028.00, which as noted in Section 5.1.1 shows a sharp hydrogen emission that does not originate from the system itself but from a background source in the galactic plane, is double-lined at the lower resolution H in two exposures and is highly asymmetric at H after masking the hydrogen cloud emission.
+Our observations of this DWD binary indicate a very large RV km s-1, which is the largest velocity difference detected amongst the full observed sample and the system is composed of two low-mass WDs of and , giving rise to a maximum orbital period of 0.4 d.
+It is possible that there are more compact systems than these in the observed sample based on our random phase sampling, and phase-resolved spectroscopy is the means to find out.
 The new double-lined systems discovered seemingly show no preference for the cooler star being the more massive of the two.
@@
 [kK] | [dex] | [] | [kK] | [dex] | [] | [] | [pc] | # | [km s-1] | [d] | ||
-J062538.73162132.02 | 14.1 | 8.03 | 0.63 | 11.3 | 7.81 | 0.50 | 1.13 | 149.5 | 1 | 7.4 | - | |
-J211927.07030002.38 | 9.4 | 7.70 | 0.44 | 5.4 | 8.52 | 0.92 | 1.36 | 93.7 | 3 | 0.6 | - | |
-J225123.02+293944.49 | 5.7 | 7.96 | 0.56 | 5.6 | 8.52 | 0.92 | 1.49 | 20.0 | 3 | 5.4 | - |
+J062538.73162132.02 | 14.0 | 7.99 | 0.60 | 11.3 | 7.87 | 0.53 | 1.14 | 149.8 | 1 | 6.4 | - | |
+J211927.07030002.38 | 11.1 | 8.31 | 0.80 | 6.5 | 7.56 | 0.40 | 1.19 | 95.0 | 3 | 4.1 | - | |
+J225123.02+293944.49 | 6.3 | 8.56 | 0.95 | 5.1 | 7.79 | 0.46 | 1.41 | 19.9 | 3 | 29.0 | - |
 WDJ name | T | g1 | M1 | T | g2 | M2 | MT | D | Exp | RV | Ref |
 [kK] | [dex] | [] | [kK] | [dex] | [] | [] | [pc] | # | [km s-1] | ||
-J053316.88075049.72 | 11.2 | 7.27 | 0.36 | 7.5 | 7.05 | 0.25 | 0.61 | 80.5 | 1 | - | |
-J141354.17+153020.71 | 15.3 | 7.78 | 0.50 | 13.2 | 7.97 | 0.59 | 1.09 | 147.7 | 2 | - | |
-J145011.93191408.67 | 8.2 | 7.85 | 0.51 | 6.2 | 7.70 | 0.43 | 0.94 | 48.6 | 5 | 1 | |
-J180600.36002720.92 | 24.1 | 7.60 | 0.45 | 8.5 | 7.46 | 0.37 | 0.82 | 147.7 | 3 | - | |
-J193642.54054744.38 | 10.6 | 7.99 | 0.59 | 7.9 | 7.83 | 0.50 | 1.09 | 105.3 | 3 | - | |
+J053316.88075049.72 | 11.1 | 7.22 | 0.32 | 7.7 | 7.16 | 0.29 | 0.61 | 80.5 | 1 | - | |
+J141354.17+153020.71 | 15.2 | 7.61 | 0.42 | 12.1 | 8.36 | 0.83 | 1.25 | 145.6 | 2 | - | |
+J145011.93191408.67 | 8.1 | 7.66 | 0.42 | 5.3 | 7.78 | 0.46 | 0.87 | 48.6 | 5 | 1 | |
+J180600.36002720.92 | 25.0 | 7.60 | 0.45 | 11.2 | 7.94 | 0.57 | 1.03 | 150.4 | 3 | - | |
+J193642.54054744.38 | 10.2 | 7.83 | 0.51 | 7.1 | 7.98 | 0.58 | 1.09 | 102.2 | 3 | - | |
 J204922.92+262517.50 | 16.9 | 7.83 | 0.53 | 14.6 | 8.06 | 0.65 | 1.17 | 141.7 | 4 | - | |
-J211345.93+262133.27 | 8.7 | 7.76 | 0.47 | 5.6 | 7.88 | 0.52 | 0.99 | 31.6 | 3 | 2 | |
-J234929.57+102255.57 | 14.5 | 7.80 | 0.50 | 7.0 | 7.83 | 0.50 | 1.00 | 117.3 | 3 | - |
+J211345.93+262133.27 | 8.6 | 7.79 | 0.48 | 5.8 | 7.83 | 0.49 | 0.97 | 31.4 | 3 | 2 | |
+J234929.57+102255.57 | 14.4 | 7.82 | 0.51 | 7.4 | 7.96 | 0.57 | 1.09 | 116.1 | 3 | - |
 WDJ name | T | g | T | g | T | g | T | D | Exp | p | Ref |
 [kK] | [dex] | [kK] | [dex] | [kK] | [dex] | [kK] | [pc] | # | |||
-J001353.60+204852.65 | 17.8 | 7.89 | 8.4 | 7.82 | 16.7 | 7.80 | 6.3 | 294.4 | 3 | -0.6 | - |
-J003045.78+273333.36 | 16.2 | 7.90 | 7.9 | 7.74 | 14.7 | 7.74 | * | 164.0 | 4 | -0.2 | - |
-J014511.23+313243.56 | 10.0 | 7.97 | 6.2 | 8.00 | 10.0 | 7.99 | 6.5 | 36.5 | 1 | - | 1 |
-J023117.04+285939.88 | 7.5 | 8.01 | 5.8 | 7.99 | 8.0 | 8.13 | * | 27.9 | 5 | -0.0 | - |
-J024323.67143957.37 | 9.0 | 7.63 | 5.0 | 7.86 | 9.8 | 7.92 | * | 79.4 | 3 | -99.0 | - |
-J080739.31+132110.65 | 11.7 | 7.66 | 6.8 | 7.71 | 11.0 | 7.57 | * | 114.7 | 1 | - | - |
-J084417.70+750008.79 | 14.6 | 8.07 | 6.9 | 8.00 | 13.4 | 7.95 | 6.3 | 101.7 | 4 | -0.0 | - |
-J084634.41+194505.18 | 10.9 | 7.61 | 7.5 | 7.59 | 10.4 | 7.51 | * | 147.7 | 1 | - | - |
-J102459.83+044610.50 | 12.3 | 7.61 | 5.7 | 7.66 | 12.4 | 7.61 | * | 43.2 | 1 | - | - |
-J104709.19+345346.65 | 9.1 | 7.63 | 5.2 | 7.82 | 9.1 | 7.69 | * | 79.3 | 1 | - | - |
-J113100.20+493826.27 | 10.9 | 7.75 | 6.9 | 8.03 | 10.8 | 7.71 | 6.5 | 98.7 | 1 | - | - |
-J113347.81+624313.29 | 6.9 | 7.65 | 5.6 | 8.88 | 6.4 | 7.89 | 6.0 | 45.8 | 1 | - | 2 |
-J131257.90+580511.29 | 11.6 | 8.01 | 6.4 | 8.05 | 11.4 | 8.00 | 6.4 | 30.9 | 2 | -0.0 | - |
-J135738.69025819.41 | 16.9 | 7.75 | 14.1 | 8.23 | 15.7 | 7.54 | 6.6 | 136.9 | 2 | -0.2 | - |
-J154214.21034131.29 | 11.6 | 8.16 | 8.1 | 7.83 | 11.0 | 7.95 | 5.3 | 53.1 | 1 | - | - |
-J172740.51+102337.94 | - | - | - | - | 10.1 | 8.04 | 7.8 | 58.9 | 5 | -9.9 | - |
-J191329.91+163822.12 | 8.4 | 7.94 | 6.7 | 7.91 | 8.9 | 8.03 | * | 77.0 | 5 | -1.0 | - |
-J192359.24+214103.62 | 9.5 | 7.77 | 5.7 | 8.08 | 10.0 | 8.00 | 6.4 | 35.1 | 3 | -0.0 | - |
-J193833.62092519.87 | 16.0 | 8.12 | 8.4 | 7.68 | 12.5 | 7.68 | 6.9 | 115.2 | 3 | -6.7 | - |
-J204517.85+810503.40 | 8.6 | 8.16 | 7.6 | 7.97 | 9.3 | 8.12 | * | 67.0 | 3 | -1.7 | 1 |
-J205650.56+062149.68 | 10.4 | 7.83 | 6.5 | 7.95 | 10.3 | 7.83 | 6.4 | 95.3 | 3 | -0.6 | - |
-J213616.39131834.50 | 10.2 | 7.83 | 5.1 | 7.92 | - | - | - | 23.4 | 3 | -0.1 | - |
-J214632.37+155039.08 | 8.6 | 7.89 | 6.8 | 7.96 | 9.1 | 7.93 | * | 72.8 | 3 | -0.2 | - |
-J221052.87045540.80 | 7.9 | 7.86 | 6.8 | 7.87 | - | - | - | 63.4 | 3 | -0.1 | - |
-J231443.05073417.85 | 14.1 | 8.10 | 9.6 | 7.93 | - | - | - | 75.2 | 3 | -0.1 | - |
-J231519.82052900.27 | 6.6 | 7.69 | 5.7 | 7.71 | 6.5 | 7.63 | * | 54.7 | 3 | -2.3 | - |
-J233041.67+110206.43 | 21.5 | 7.96 | 6.5 | 8.07 | 21.4 | 7.96 | 6.4 | 98.8 | 1 | - | - |
+J001353.60+204852.65 | 17.7 | 7.90 | 8.3 | 7.83 | 16.8 | 7.78 | 7.2 | 291.0 | 3 | -0.6 | - |
+J003045.78+273333.36 | 16.1 | 7.86 | 7.6 | 7.74 | 14.6 | 7.70 | * | 165.1 | 4 | -0.2 | - |
+J014511.23+313243.56 | 9.6 | 7.87 | 6.4 | 8.00 | 9.6 | 7.86 | 6.5 | 36.7 | 1 | - | 1 |
+J023117.04+285939.88 | 7.5 | 7.96 | 5.8 | 8.08 | 7.5 | 7.82 | * | 27.9 | 5 | -0.0 | - |
+J024323.67143957.37 | 9.2 | 7.72 | 5.1 | 7.61 | 9.1 | 7.70 | * | 80.2 | 3 | -99.0 | - |
+J080739.31+132110.65 | 11.6 | 7.68 | 7.3 | 7.80 | 11.1 | 7.61 | * | 115.0 | 1 | - | - |
+J084417.70+750008.79 | 14.5 | 8.04 | 7.0 | 8.01 | 13.3 | 7.96 | 7.1 | 102.4 | 4 | -0.0 | - |
+J084634.41+194505.18 | 10.9 | 7.65 | 7.7 | 7.61 | 10.1 | 7.44 | * | 146.5 | 1 | - | - |
+J102459.83+044610.50 | 12.4 | 7.62 | 5.7 | 7.64 | 12.1 | 7.60 | * | 43.1 | 1 | - | - |
+J104709.19+345346.65 | 9.2 | 7.70 | 5.4 | 7.69 | 9.0 | 7.63 | * | 79.4 | 1 | - | - |
+J113100.20+493826.27 | 11.6 | 7.91 | 6.9 | 7.71 | 10.7 | 7.72 | 7.3 | 99.3 | 1 | - | - |
+J113347.81+624313.29 | - | - | - | - | 6.4 | 7.92 | 6.0 | 45.8 | 1 | - | 2 |
+J131257.90+580511.29 | 11.6 | 8.01 | 6.7 | 8.12 | 11.2 | 7.93 | 6.4 | 30.9 | 2 | -0.0 | - |
+J135738.69025819.41 | 17.1 | 7.92 | 14.8 | 7.98 | 16.0 | 7.57 | 6.9 | 137.2 | 2 | -0.2 | - |
+J154214.21034131.29 | 11.5 | 8.07 | 7.9 | 8.07 | 10.8 | 7.89 | 5.5 | 51.9 | 1 | - | - |
+J172740.51+102337.94 | - | - | - | - | 10.5 | 8.23 | 7.8 | 59.0 | 5 | -9.9 | - |
+J191329.91+163822.12 | 8.3 | 7.88 | 6.7 | 8.03 | 8.4 | 7.78 | * | 77.0 | 5 | -1.0 | - |
+J192359.24+214103.62 | 9.4 | 7.81 | 6.1 | 7.81 | 9.5 | 7.80 | 6.4 | 35.3 | 3 | -0.0 | - |
+J193833.62092519.87 | 15.9 | 8.08 | 8.0 | 7.66 | 12.5 | 7.73 | 8.7 | 114.8 | 3 | -6.7 | - |
+J204517.85+810503.40 | 8.7 | 8.19 | 7.8 | 8.01 | 8.9 | 7.93 | * | 67.0 | 3 | -1.7 | 1 |
+J205650.56+062149.68 | 11.0 | 7.98 | 6.5 | 7.67 | 10.1 | 7.77 | 6.7 | 95.0 | 3 | -0.6 | - |
+J213616.39131834.50 | 10.3 | 7.85 | 5.0 | 8.03 | - | - | - | 23.4 | 3 | -0.1 | - |
+J214632.37+155039.08 | 8.5 | 7.83 | 6.9 | 7.96 | 8.6 | 7.79 | * | 74.5 | 3 | -0.2 | - |
+J221052.87045540.80 | 7.9 | 7.84 | 6.9 | 7.90 | - | - | - | 63.5 | 3 | -0.1 | - |
+J231443.05073417.85 | 14.0 | 7.99 | 9.5 | 8.10 | - | - | - | 75.1 | 3 | -0.1 | - |
+J231519.82052900.27 | 6.5 | 7.65 | 5.9 | 7.88 | 6.3 | 7.48 | * | 54.4 | 3 | -2.3 | - |
+J233041.67+110206.43 | 21.7 | 7.97 | 7.0 | 8.10 | 21.5 | 7.94 | 6.4 | 99.2 | 1 | - | - |
 WDJ name | T | g | M | D | Spec | p | Ref | Other Name |